The force between two short electric dipoles seperated by a distance r is directly proportional to
a r^2
b r^4
c 1/r^2
d 1/r^4

Grade:12

2 Answers

Aseespal
6 years ago
The force between two dipoles i
Abhishek Rai
6 years ago
I tried to find the formula for this, but could only find the force on dipoles in uniform fields. Of course, in that case, if the dipoles are in-line, the force is zero. However, the field from a dipole is not uniform an is E = K*p/r³, where p = dipole moment. I will do my best to derive a result. The force on the first charge is E*q = K*p*q/r³. If the second charge is a distance d from the first, the field there is K*p/(r + d)³ and the force is -K*p*q/(r + d)³ The net force on the dipole is then K*p*q/r³ - K*p/(r + d)³ K*p*q*[1/r³ - 1/(r + d)³)] K*p*q*[(r + d)³ - r³]/[r³*(r + d)³] Divide num & denom by r³: K*p*q*[(1 + d/r)³ - 1]/(r + d)³ Expand the binomial in the numerator: K*p*q*[1 + 3*d/r + 3(d/r)² + (d/r)³ - 1]/(r + d)³ For d